# Does rare, noise-induced, bypass transition in plane Couette flow bypass instantons ?

### v1

@@ 5.3 Statistical description of the paths
    Finally, a peculiar situation has been reported following the computation of the instanton going from a non linear traveling wave to the laminar flow in two dimensional plane Poiseuille flow [Wan et al., 2015]. In that case, sketched in figure 15 (b), the instanton reaches the neighbourhood of the separatrix near a flow configuration termed a mediator state, travels along the separatrix toward the unstable fixed point then relaxes toward its destination. The peculiarity of that situation is that once the flow has reached the mediator point, a small numerical error in instanton calculation, or a small amount of noise in a stochastic system, can push the dynamics on the other side of the separatrix : the dynamics then travels near deterministically toward the final state $\mathcal{B}$. Let us term $\mathcal{I}_{\mathcal{A}\rightarrow\mathcal{B}}$ the full action of the instanton going from $\mathcal{A}$ to $\mathcal{B}$, $\mathcal{I}_{\mathcal{A}\rightarrow m}$, the action of a trajectory following the instanton from $\mathcal{A}$ to the mediator state, $\mathcal{I}_{m\rightarrow s}$ the action of a trajectory following the instanton from the mediator state to the saddle and $\mathcal{I}_{m\rightsquigarrow\mathcal{B}}$ the action of a trajectory starting on the mediator and being slightly perturbed to cross the separatrix and then travel deterministically toward $\mathcal{B}$. The action of deterministic paths is always zero: this means that in the limit of noise variance going to zero, the action of any path starting on the $\mathcal{B}$ side of the separatrix to $\mathcal{B}$ that has any chance of being observed is zero, because it will follow a deterministic path. This means that $\mathcal{I}_{\mathcal{A}\rightarrow\mathcal{B}}=\mathcal{I}_{\mathcal{A}\rightarrow m}+\mathcal{I}_{m\rightarrow s}\simeq\mathcal{I}_{\mathcal{A}\rightarrow m}$: the action of the instanton is that of traveling toward the mediator point then to the saddle and then relaxing. Moreover the part of the path along the separatrix has a negligible action $\mathcal{I}_{m\rightarrow s}\ll\mathcal{I}_{\mathcal{A}\rightarrow m}$ because it is a near deterministic path in the stable manifold of the saddle. Similarly, the action of a path traveling toward the mediator along the instanton then receiving a small amount of noise to cross the separatrix there (as sketched in blue sketched in figure 15 (b)) is $\mathcal{I}_{\mathcal{A}\rightsquigarrow\mathcal{B}}=\mathcal{I}_{\mathcal{A}\rightarrow m}+\mathcal{I}_{m\rightsquigarrow\mathcal{B}}\simeq\mathcal{I}_{\mathcal{A}\rightarrow m}$. The action $\mathcal{I}_{\mathcal{A}\rightarrow\mathcal{B}}$ and $\mathcal{I}_{\mathcal{A}\rightsquigarrow\mathcal{B}}$ are very similar and the corresponding probabilities and mean first passage time respectively go to zero and to infinity exponentially as very similar rates. At finite though small noise variance, the probability of observing a trajectory going along the instanton all the way becomes small compared to observing a trajectory going toward the mediator then crossing toward $\mathcal{B}$ there, because the instanton requires the trajectory not to be perturbed once the mediator has been reached. Therefore, computations by means of numerical simulation of the stochastic process (like DNS or AMS) would give the trajectory crossing near the mediator. Moreover, trajectories would still concentrate up to the mediator. However, the estimate of the probability and the mean first passage time would go like $\exp(\pm\beta\mathcal{I}_{\mathcal{A}\rightarrow m})$ and thus be exponential and very close to the one predicted by the Freidlin Wentzel principle of large deviations. In our case, the exponential dependence of $\alpha$ and $T$, the crossing of the separatrix away from the saddle and the concentration of paths are consistent with this scenario. We note however that the computed path do not concentrate on top of one another in planes computed using kinetic energy. This could be an effect of the finite noise variance in these quadratic spatially averaged quantities. Indeed, the mechanical scenario is not changed from one value of $\beta$ to another.
@@
As far as we can deduce from the data presented here, scenario iv is the most similar to build up in plane Couette flow. If it is the case, this would mean that we would observe an actual instanton all along the trajectory, not just until the mediator or turning point only, for exceedingly small values of the noise. Moreover, if we follow this line of reasoning, we cannot rule out the possibility that there could be a mediator even closer to the laminar state, comparable to a minimal seed, that would only be reached at minuscule values of noise variance. It could be even beyond the reach of rare events simulations methods. The case of build up in plane Couette flow thus helped us identifying a different regime of convergence toward instantons, already pointed out in two dimensional plane Poiseuille flow [Wan et al., 2015] and can be linked to observations made in other systems where crossing away from the saddle and very flat dependence of the action on the crossing point in model of oceanic flows [Börner et al., 2023]. This behaviour could thus manifest itself in more complex bistable geophysical flows.
@@
Figure 15: (a) Sketch of the two metastable states $\mathcal{A}$ and $\mathcal{B}$ in phase space, separated by a separatrix (green curve), on which we have the transition state: the most accessible saddle point (red arrows indicating stable and unstable directions). In the limit of vanishing energy injection rate $\epsilon\rightarrow 0$, the reactive trajectory should follow an instanton (in cyan): on this first sketch, we represent an instanton that nears the separatrix only as it reaches the saddle along mostly unstable directions. For small but finite energy injection rate $\epsilon$, we would expect that trajectories concentrate around the instanton (in yellow). Instead, we systematically observe trajectories that concentrate around a path that crosses the separatrix away from the instanton and the most accessible fixed point (in blue). For these two types of paths: the full line indicates the average trajectory, while the dashed lines indicate where the zone that contains the majority of paths. (b) Modification of sketch (a) in the case where the instanton approaches the separatrix near a mediator state, then goes along the separatrix approaches the nearest fixed point and finally travels towards $\mathcal{B}$, with a scenario similar to the one described by [Wan et al., 2015].
@@ Acknowledgment
The author thanks Institut PPrime, ENSMA and their computational facilities, where this work was started, for its hospitality. The author thanks the help of the computational resources of the PSMN platform of ENS de Lyon, as well as the hospitality of the laboratoire de Physique de l’ENS de Lyon, where part of the computation of this work have been performed. Additional computational resources were granted by Mésocentre de Calcul Scientifique Intensif of the University of Lille and this work was granted access to the HPC resources of IDRIS under the allocation 2022-A0122A01741 made by GENCI. The author also thanks the support from CICADA (centre de calcul interactifs) (now mésocentre Azzurra) of University Nice Sophia-Antipolis where some early computations have been performed. The author finally thanks F. Bouchet, R. Börner, J.-C. Vassilicos and J.-P. Laval for inspiring discussions.
@@ Appendix A Structure of the forcing noise
Appendix B Parameter used in definition of reaction coordinate in each set of simulations
@@
In section 2.2.2, we presented the set up of the reaction coordinates (Eq. (11)), using three free parameters $E_{\mathcal{A}}$ to define the laminar state $\mathcal{A}$, $\sigma$ to define the hypersurface $\mathcal{C}$ that surrounds it and $E_{\mathcal{B}}$ to define the turbulent state (Fig. 3 (a)). Most AMS computations used a systematic definition of these three parameters with $E_{\mathcal{B}}=0.06$, and the parameters controlling $E_{\mathcal{A}}$ and $\sigma$ reported in table 1
@@
Appendix C Convergence of AMS computations
In this appendix, in complement to the precision tests of section 4.1 (Fig. 13 (d)), we will check that the number of clones used in AMS computation was sufficient. For this matter, we present convergence results with the number of clones for the two types of forcings in domains of size $L_{x}\times L_{z}=6\times 4$ at Reynolds number $R=500$ using standard resolution $N_{x}=32$, $N_{y}=27$, $N_{z}=24$ and $\Delta t=0.05$. Isotropic forcing was added using $\beta=4\cdot 10^{5}$ and $\beta=5\cdot 10^{5}$ with $N_{r}=6$, and $L_{c}=1.0$. Divergence free forcing was added using $\beta=2.5\cdot 10^{5}$ and $3\cdot 10^{5}$ using $L_{c}=1$ and $N_{r}=4$. In all four cases, the crossing probabilities are not that small, so that the reactive trajectories and their properties can be computed by DNS with a high enough degree of precision.
@@ Appendix C Convergence of AMS computations
where $\langle\kappa\rangle_{o}$ is the average number of iterations over AMS runs and $\langle r\rangle_{o}$ is the average proportion of clones reaching $\mathcal{B}$ over AMS runs. It has been demonstrated that this quantity is larger than or equal to one, and it is one when the committor is used as reaction coordinate [Cérou et al., 2019]. In parallel, it has also been shown that the closer a reaction coordinate is to the committor, the most precise the AMS estimates are when a finite number of runs are used. As a consequence, this quantity is used as an estimate of the quality of the computation because the better the reaction coordinate, the smaller is $\sigma$. A very large $\sigma$ would be that the estimates of $\alpha$ and $T$ are incorrect, or even that the paths are incorrectly selected [Rolland and Simonnet, 2015, Lucente et al., 2022]. In that case what is typically observed in the selection of shorter though less probable paths. In the two test cases considered here (Fig. 16 (c)), we can see that $\sigma$ is slightly larger than one. This ensures correct estimates, but also tells us that there is room for improvement of the reaction coordinate.
Refer to caption Refer to caption Refer to caption
Figure 16: In a domain of size $L_{x}\times L_{z}=6\times 4$ at Reynolds number $R=500$ for four different forcings (see § C for details). (a) Crossing $\alpha$ probability as a function of the number of clones, computed by AMS or DNS. The error bars correspond to the $66\%$ confidence interval of AMS computations, the dotted line corresponds to the 66% interval of confidence of DNS computations. (b) normalised standard deviation (Eq. (35)) of the crossing probability $\alpha$ as a function of the number of clones, compared to $1$. (c) Average duration $\tau$ of reactive trajectories as a function of the number of clones, computed by means of AMS and DNS for the same cases and the same line codes as (a).

### v2

@@ 5.3 Statistical description of the paths
    Finally, a peculiar situation has been reported following the computation of the instanton going from a non linear traveling wave to the laminar flow in two dimensional plane Poiseuille flow Wan et al., (2015). In that case, sketched in figure 17 (b), the instanton reaches the neighbourhood of the separatrix near a flow configuration termed a mediator state, travels along the separatrix toward the unstable fixed point then relaxes toward its destination. The peculiarity of that situation is that once the flow has reached the mediator point, a small numerical error in instanton calculation, or a small amount of noise in a stochastic system, can push the dynamics on the other side of the separatrix : the dynamics then travels near deterministically toward the final state $\mathcal{B}$. Let us term $\mathcal{I}_{\mathcal{A}\rightarrow\mathcal{B}}$ the full action of the instanton going from $\mathcal{A}$ to $\mathcal{B}$, $\mathcal{I}_{\mathcal{A}\rightarrow m}$, the action of a trajectory following the instanton from $\mathcal{A}$ to the mediator state, $\mathcal{I}_{m\rightarrow s}$ the action of a trajectory following the instanton from the mediator state to the saddle and $\mathcal{I}_{m\rightsquigarrow\mathcal{B}}$ the action of a trajectory starting on the mediator and being slightly perturbed to cross the separatrix and then travel deterministically toward $\mathcal{B}$. The action of deterministic paths is always zero: this means that in the limit of noise variance going to zero, the action of any path starting on the $\mathcal{B}$ side of the separatrix to $\mathcal{B}$ that has any chance of being observed is zero, because it will follow a deterministic path. This means that $\mathcal{I}_{\mathcal{A}\rightarrow\mathcal{B}}=\mathcal{I}_{\mathcal{A}\rightarrow m}+\mathcal{I}_{m\rightarrow s}\simeq\mathcal{I}_{\mathcal{A}\rightarrow m}$: the action of the instanton is that of traveling toward the mediator point then to the saddle and then relaxing. Moreover the part of the path along the separatrix has a negligible action $\mathcal{I}_{m\rightarrow s}\ll\mathcal{I}_{\mathcal{A}\rightarrow m}$ because it is a near deterministic path in the stable manifold of the saddle. Similarly, the action of a path traveling toward the mediator along the instanton then receiving a small amount of noise to cross the separatrix there (as sketched in blue sketched in figure 17 (b)) is $\mathcal{I}_{\mathcal{A}\rightsquigarrow\mathcal{B}}=\mathcal{I}_{\mathcal{A}\rightarrow m}+\mathcal{I}_{m\rightsquigarrow\mathcal{B}}\simeq\mathcal{I}_{\mathcal{A}\rightarrow m}$. The action $\mathcal{I}_{\mathcal{A}\rightarrow\mathcal{B}}$ and $\mathcal{I}_{\mathcal{A}\rightsquigarrow\mathcal{B}}$ are very similar and the corresponding probabilities and mean first passage time respectively go to zero and to infinity exponentially as very similar rates. At finite though small noise variance, the probability of observing a trajectory going along the instanton all the way becomes small compared to observing a trajectory going toward the mediator then crossing toward $\mathcal{B}$ there, because the instanton requires the trajectory not to be perturbed once the mediator has been reached. Therefore, computations by means of numerical simulation of the stochastic process (like DNS or AMS) would give the trajectory crossing near the mediator. Moreover, trajectories would still concentrate up to the mediator. However, the estimate of the probability and the mean first passage time would go like $\exp(\pm\beta\mathcal{I}_{\mathcal{A}\rightarrow m})$ and thus be exponential and very close to the one predicted by the Freidlin Wentzel principle of large deviations. In our case, the exponential dependence of $\alpha$ and $T$, the crossing of the separatrix away from the saddle and the concentration of paths are consistent with this scenario. We note however that the computed path do not concentrate on top of one another in planes computed using kinetic energy. This could be an effect of the finite noise variance in these quadratic spatially averaged quantities. Indeed, the mechanical scenario is not changed from one value of $\beta$ to another.
@@
As far as we can deduce from the data presented here, scenario iv is the most similar to build up in plane Couette flow. If it is the case, this would mean that we would observe an actual instanton all along the trajectory, not just until the mediator or turning point only, for exceedingly small values of the noise. Moreover, if we follow this line of reasoning, we cannot rule out the possibility that there could be a mediator even closer to the laminar state, comparable to a minimal seed, that would only be reached at minuscule values of noise variance. It could be even beyond the reach of rare events simulations methods. The case of build up in plane Couette flow thus helped us identifying a different regime of convergence toward instantons, already pointed out in two dimensional plane Poiseuille flow Wan et al., (2015) and can be linked to observations made in other systems where crossing away from the saddle and very flat dependence of the action on the crossing point in model of oceanic flows Börner et al., (2023). This behaviour could thus manifest itself in more complex bistable geophysical flows. A similar behaviour was observed in a model representing a first order phase transition Grafke and Vanden-Eijnden, (2017), indicating that this type of analysis is also relevant for statistical physics.
@@
Figure 17: (a) Sketch of the two metastable states $\mathcal{A}$ and $\mathcal{B}$ in phase space, separated by a separatrix (green curve), on which we have the transition state: the most accessible saddle point (red arrows indicating stable and unstable directions). In the limit of vanishing energy injection rate $\epsilon\rightarrow 0$, the reactive trajectory should follow an instanton (in cyan): on this first sketch, we represent an instanton that nears the separatrix only as it reaches the saddle along mostly unstable directions. For small but finite energy injection rate $\epsilon$, we would expect that trajectories concentrate around the instanton (in yellow). Instead, we systematically observe trajectories that concentrate around a path that crosses the separatrix away from the instanton and the most accessible fixed point (in blue). For these two types of paths: the full line indicates the average trajectory, while the dashed lines indicate where the zone that contains the majority of paths. (b) Modification of sketch (a) in the case where the instanton approaches the separatrix near a mediator state, then goes along the separatrix approaches the nearest fixed point and finally travels towards $\mathcal{B}$, with a scenario similar to the one described by Wan et al., (2015).
@@ Acknowledgment
The author thanks Institut PPrime, ENSMA and their computational facilities, where this work was started, for its hospitality. The author thanks the help of the computational resources of the PSMN platform of ENS de Lyon, as well as the hospitality of the laboratoire de Physique de l’ENS de Lyon, where part of the computation of this work have been performed. Additional computational resources were granted by Mésocentre de Calcul Scientifique Intensif of the University of Lille and this work was granted access to the HPC resources of IDRIS under the allocation 2022-A0122A01741 made by GENCI. The author also thanks the support from CICADA (centre de calcul interactifs) (now mésocentre Azzurra) of University Nice Sophia-Antipolis where some early computations have been performed. The author finally thanks F. Bouchet, R. Börner, J.-C. Vassilicos and J.-P. Laval for inspiring discussions, as well as E. Vanden–Eijnden for pointing interesting references.
@@ Appendix A Structure of the forcing noise
Appendix B Derivation of the kinetic energy budget
In section 2.1.2, we have presented the ensemble averaged budget of spatially averaged kinetic energy (Eq. (7)) which resembles the Reynolds–Orr energy equation, with the addition of an energy injection rate (Eq. (8)). We have stated succinctly that this budget, and the similar budget of total energy (Eq. (10)), was obtained by application of the Itô Lemma for change of variables in stochastic differential equations. In this appendix, we present the main steps to go from the Stochastically forced Navier–Stokes equations (2) to the energy budgets (Eq. (7), (10)).
We first present the principle of such a change of variable for a scalar time dependent variable $f(t)$ following a time discretised Langevin equation of the type
$f(t+{\rm d}t)-f(t)=df=F(f(t)){\rm d}t+B(f(t)){\rm d}W\,,$ (18)
where ${\rm d}W(t)$ is a Wiener process discretised with Itô convention. This means that $\langle{\rm d}W(t)\rangle=0$, $\langle{\rm d}W(t){\rm d}W(t^{\prime})\rangle={\rm d}t\delta(t-t^{\prime})$, and the noise covariance $B(f)$ and the Wiener process are uncorrelated $\langle B(f(t)){\rm d}W(t)\rangle=0$. The Itô Lemma (see (Gardiner,, 2009) § 4) states how a change of variable towards the Langevin equation for $g(f)$ is performed. Its application leads to a stochastic differential equation where the noise is still interpreted with an Itô convention. It shows that as soon as $g$ is non linear, an additional deterministic term should be included, on top of the operations that would be naturally performed to change variables in an ordinary differential equation. The resulting Langevin equation reads
$dg=\frac{dg}{df}F(f){\rm d}t+\frac{1}{2}B^{2}\frac{d^{2}g}{df^{2}}{\rm d}t+\frac{dg}{df}B(f){\rm d}W\,.$ (19)
In this stochastic differential equation, the terms $\frac{dg}{df}F(f){\rm d}t$ and $\frac{dg}{df}B(f){\rm d}W$ are what would arise from a change of variable in an ordinary differential equation, and can be obtained by any standard mean. The additional term $\frac{1}{2}B^{2}\frac{d^{2}g}{df^{2}}{\rm d}t$, sometimes called a ghost drift, accounts for the fact that the correlation between the noise and the new variable, while physically existent, are excluded from the noise term by the Itô convention in equation (19). This convention leads to $\left\langle\frac{dg}{df}B(x){\rm d}W\right\rangle=0$ when equation (19) is ensemble averaged, to give
$\left\langle\frac{dg}{dt}\right\rangle=\left\langle\frac{dg}{df}F\right\rangle+\left\langle\frac{1}{2}B^{2}\frac{d^{2}g}{df^{2}}\right\rangle\,.$ (20)
This new ghost drift term is absent if $g$ is affine in $f$, and it is an additive constant if $g$ is quadratic in $f$ and the noise covariance $B$ is independent on the variable $f$. This last situation represents a one degree of freedom version of our derivation. We now present the operation in the in more general cases of multidimensional Langevin equations for vector function $\vec{f}(t)$ with $N$ components. In that case, $B$ is an $N\times N$ matrix, applied to a vector of independent Wiener processes. Itô Lemma then tells us that to perform a change of variables $g\left(\vec{f}\right)$, the first part of the resulting equation can be obtained through the classical means, and that we need to find a convenient mean to calculate the resulting ghost drift that can formally be written as
$\frac{1}{2}\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{N}B_{ik}B_{kj}\frac{\partial^{2}g}{\partial f_{i}\partial f_{j}}\,,$ (21)
where $BB^{\dagger}=\sum_{k=1}^{N}B_{ik}B_{kj}$ is the covariance matrix of the multidimensional noise, which is contracted against the Hessian matrix of $g$.
In our case, we change variables from the velocity fields $\mathbf{u}$ to the spatially averaged kinetic energy. This change of variables is quadratic in the variable $\mathbf{u}$, so that we expect a purely additive constant ghost drift. In order to calculate this drift, we will use two properties:
  1. i)
    the noise covariance is easily written in Fourier–Chebyshev space (Eq. (3), § A), the vector $\vec{f}$ that we manipulate then consists of all the Fourier and Chebyshev modes of all the components of $\mathbf{u}$.
  2. ii)
    the spatially averaged kinetic energy can also be written simply in Fourier–Chebyshev space using Parseval Theorem
    $E_{k}(t)=\sum_{\begin{subarray}{c}n_{x}=\frac{N_{x}}{2}-1\\ n_{y}=0\\ n_{z}=\frac{N_{z}}{2}-1\end{subarray}}^{\begin{subarray}{c}n_{z}=\frac{N_{z}}{2}+1\\ n_{y}=N_{y}-1\\ nx=\frac{N_{x}}{2}+1\end{subarray}}\frac{\left(|\hat{u}_{x}|^{2}+|\hat{u}_{y}|^{2}+|\hat{u}_{z}|^{2}\right)(n_{x},n_{y},n_{z})}{2}\,$ (22)
    because $\mathbf{u}$ is written in the orthonormal basis of Fourier and Chebyshev modes in our discretisation. Moreover, using this spectral representation makes the description of the variable involved in the Langevin equation as a vector with components labelled by integer much simpler, even in the limit of continuous fields $N_{x},N_{y},N_{z}\rightarrow\infty$. We can still use derivative with respect to the modes instead of differentials in the Hessian. In that basis, the Hessian of the change of variables is then the identity matrix (the $\frac{1}{2}$ factor and the factor $2$ coming from the derivative cancel out).
This means that the contraction (Eq. (21)) of the Hessian of the change of variables with the covariance matrix of the noise in spectral space is just its trace, leading to the energy injection rate of equation (8). Note that there is an additional simplification between the factor $\frac{2}{\beta}$ in the noise covariance and the factor $\frac{1}{2}$ in front of the ghost drift.
The same procedure is applied to obtain equation (10). This leads to the same additive constant as a ghost drift because the difference in the changes of variables (from $\mathbf{u}$ to $E_{k}$ (Eq. (4)) and from $\mathbf{u}$ to $E_{\rm tot}$, (Eq. (10))) only lies in constant or linear terms in components of $\mathbf{u}$: these differences do not leave a trace in the Hessian of the change of variables. Meanwhile, after an ensemble average that cancels out the noise, the rest of the differential equation is the same as in the deterministic case.
@@ Appendix B Derivation of the kinetic energy budget
Appendix C Parameter used in definition of reaction coordinate in each set of simulations
@@
In section 2.2.2, we presented the set up of the reaction coordinates (Eq. (11)), using three free parameters $E_{\mathcal{A}}$ to define the laminar state $\mathcal{A}$, $\sigma$ to define the hypersurface $\mathcal{C}$ that surrounds it and $E_{\mathcal{B}}$ to define the turbulent state (Fig. 3 (a)). Most AMS computations used a systematic definition of these three parameters with $E_{\mathcal{B}}=0.06$, and the parameters controlling $E_{\mathcal{A}}$ and $\sigma$ reported in table 4.
@@
Appendix D Convergence of AMS computations
In this appendix, in complement to the precision tests of section 4.1 (Fig. 14 (c)), we will check that the number of clones used in AMS computation was sufficient. For this matter, we present convergence results with the number of clones for the two types of forcings in domains of size $L_{x}\times L_{z}=6\times 4$ at Reynolds number $R=500$ using standard resolution $N_{x}=32$, $N_{y}=27$, $N_{z}=24$ and $\Delta t=0.05$. Isotropic forcing was added using $\beta=4\cdot 10^{5}$ and $\beta=5\cdot 10^{5}$ with $N_{r}=6$, and $L_{c}=1.0$. Divergence free forcing was added using $\beta=2.5\cdot 10^{5}$ and $3\cdot 10^{5}$ using $L_{c}=1$ and $N_{r}=4$. In all four cases, the crossing probabilities are not that small, so that the reactive trajectories and their properties can be computed by DNS with a high enough degree of precision.
Refer to caption Refer to caption Refer to caption
Figure 18: In a domain of size $L_{x}\times L_{z}=6\times 4$ at Reynolds number $R=500$ for four different forcings (see § D for details). (a) Crossing $\alpha$ probability as a function of the number of clones, computed by AMS or DNS. The error bars correspond to the $66\%$ confidence interval of AMS computations, the dotted line corresponds to the 66% interval of confidence of DNS computations. (b) normalised standard deviation (Eq. (23)) of the crossing probability $\alpha$ as a function of the number of clones, compared to $1$. (c) Average duration $\tau$ of reactive trajectories as a function of the number of clones, computed by means of AMS and DNS for the same cases and the same line codes as (a).
@@ Appendix D Convergence of AMS computations
where $\langle\kappa\rangle_{o}$ is the average number of iterations over AMS runs and $\langle r\rangle_{o}$ is the average proportion of clones reaching $\mathcal{B}$ over AMS runs. It has been demonstrated that this quantity is larger than or equal to one, and it is one when the committor is used as reaction coordinate Cérou et al., (2019). In parallel, it has also been shown that the closer a reaction coordinate is to the committor, the most precise the AMS estimates are when a finite number of runs are used. As a consequence, this quantity is used as an estimate of the quality of the computation because the better the reaction coordinate, the smaller is $\sigma$. A very large $\sigma$ would be that the estimates of $\alpha$ and $T$ are incorrect, or even that the paths are incorrectly selected Rolland and Simonnet, (2015); Lucente et al., (2022). In that case what is typically observed in the selection of shorter though less probable paths. In the two test cases considered here (Fig. 18 (c)), we can see that $\sigma$ is slightly larger than one. This ensures correct estimates, but also tells us that there is room for improvement of the reaction coordinate.
@@
The overlap of intervals of confidences of estimates by means of AMS as well as the closeness of the rescaled variance to one indicates that AMS computations in domains of size $L_{x}\times L_{z}=6\times 4$ are performed without errors. This gives us further confidence that the reaction coordinate $\phi$ defined in section 2.2.2 correctly weighs the advance of reactive trajectories by being affine in kinetic energy. To our knowledge, nothing makes this reaction coordinate irrelevant when the size of the system is increased to $L_{x}\times L_{z}=36\times 24$: the flow transits toward turbulence by near monotonously increasing its kinetic energy, even when the flow displays spanwise localisation during the transition. Some problem would occur if the flow had to systematically decrease its kinetic energy during transition, or give some importance to a process that does not leave a trace in total kinetic energy. We expect these assertions to remain valid if AMS is used to compute reactive trajectories in larger domains, where the flow displays both spanwise and streamwise localisation during the transition. We note however that some care should be given in the definition of the thresholds defining sets $\mathcal{A}$ and $\mathcal{C}$ (Fig. 3 (a)). Indeed, with spatial localisation of velocity tubes in the flow in the early stages of the trajectory, the total kinetic energy is reduced as the area in which we find the non trivial flow is reduced. If said thresholds are too large, the early stages of AMS computation risk being too long as the algorithm is not used to generate the localised velocity tubes. Note finally that another effect can impair the quality of AMS computations: a slow separation of trajectories in phase space. While this can happen in principle for any type of physical system, this effect has mostly be seen in system not forced by any noise, and modifications of AMS have been performed to ensure that AMS computation can generate enough variability in the obtained trajectories or even function at all Rolland, (2022). No such problem has been observed in the computation of build up trajectories so far. For such a problem to manifest itself in domains larger than $L_{x}\times L_{z}=36\times 24$, where both spanwise and streamwise localisation of turbulence happens, the physics of transition would have to change drastically.